# Using Dynamic Safety Margins as Control Barrier Functions

### v2

@@ Abstract
This paper provides a systematic approach to design control barrier functions (CBFs) for arbitrary state and input constraints using tools from the reference governor literature. In particular, it is shown that dynamic safety margins (DSMs) are CBFs for an augmented system obtained by concatenating the state and a virtual reference. The proposed approach is agnostic to the relative degree and can handle multiple state and input constraints using the control-sharing property of CBFs. The construction of CBFs using Lyapunov-based DSMs is then investigated in further detail. Numerical simulations show that the method outperforms existing DSM-based approaches, while also guaranteeing safety and persistent feasibility of the associated optimization program.
@@ 1 Introduction
Given a smooth state constraint set, it is usually straightforward to design a differentiable function of the state vector that is positive for constraint-admissible states, and negative otherwise. We call such functions candidate CBFs. They are commonly used in practice to design CBF-based controllers and, with tuning, can achieve great performance (e.g., [2]). However, candidate CBFs lack many safety guarantees that valid CBFs provide. In particular, the CBF-based optimization program governing the control law may become infeasible when using candidate CBFs, which is unacceptable for safety-critical systems. Ultimately, CBFs are certificates of control invariance. Therefore, if the original state constraint set is not control invariant, it is impossible to find a valid CBF associated to it. In these cases, we must resort to finding CBFs for a, hopefully large, control invariant subset of the state constraint set. In this work, we use tools from the reference governor literature [4] to systematically design valid CBFs given arbitrary state and input constraints.
@@ 1.1 Related Works
Most relevant to the present manuscript, we review some works that study the design of CBFs and feasibility of the CBF-based optimization program. In [14], sum of squares (SOS) techniques are used to synthesize CBFs for semi-algebraic sets and polynomial systems using an alternating descent algorithm. In [15], the authors use SOS techniques to synthesize polynomial control Lyapunov functions, and then present an extension to construct CBFs. In [16], the authors introduce a variable modifying the decay rate of the CBF to guarantee feasibility of the CBF-based program when the state is in the interior of the safe set even under input constraints. However, their approach provides no feasibility guarantee when the state is on the boundary of the safe set. In [17], the authors design CBFs for Euler–Lagrange systems. While the results are promising and the class of systems is relevant, the approach is limited to box constraints. In [18], the authors learn CBFs from human demonstrations but they provide no guarantees that the learned CBFs are valid. In [19], given a valid CBF, the authors use learning to construct a CBF for a larger control invariant set using a combination of safe and unsafe trajectories of the system. Nonetheless, finding the initial, valid CBF remains an open question. In [20], the authors use a backup control policy to enlarge a small (but easy to find) control invariant set. The enlarged control invariant set is then used to derive a backup CBF. These backup CBFs rarely have a closed form, which makes them difficult to implement. In [21], the authors use maximal output admissible sets (MOASs) to design discrete-time CBFs for arbitrary state and input constraints. While the approach works well for linear systems, finding the MOAS for nonlinear systems remains an open question.
@@
This work expands the underlying theme of [20, 21], where CBFs are obtained starting from a prestabilizing (or backup) controller. To do this, we adopt the notion of dynamic safety margin (DSM) from the explicit reference governor (ERG) framework [22]. Given a fixed reference, a DSM measures the distance to constraint violation of a prestabilized system trajectory as it converges to the associated equilibrium point. In [23], the authors define governor-parameterized barrier functions (PBFs), which are based on CBFs, and show that they are DSMs. While the results are limited to linear systems, their work is one of the few connections between the reference governor and CBF literatures. In contrast to their result, we show that DSMs are CBFs for an augmented system consisting of the concatenated state and prestabilizing controller reference vectors. These works serve as motivation for exploring the connections between reference governor approaches like the command governor (CG) [24], which filters a nominal reference, and CBF-based control, which filters a nominal input. Both approaches ultimately modify a nominal, but potentially unsafe signal to guarantee closed-loop safety (see Fig. 1).
@@ 1.2 Contributions
  • We show that dynamic safety margins (DSMs) are control barrier functions (CBFs) for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM.
  • We show that the DSM-based CBFs support constraints with arbitrary relative degree.
  • We show that DSMs obtained for the same prestabilizing controller are control-sharing CBFs. That is, they can be simultaneously imposed in the CBF-based program and feasibility is retained.
  • We extend the Lyapunov-based DSMs presented in [22] to inadmissible references. This extension is needed for DSM-based CBFs.
  • We provide conditions on the DSM and the constraints that guarantee the CBF-based controller is locally Lipschitz contionuous. This guarantees uniqueness of the closed-loop trajectory.
  • We present a comparison of our proposed approach with the explicit reference governor (ERG) [22] and candidate CBFs in two nonlinear systems. The numerical simulations show that the DSM-based CBFs we propose outperform the ERG, while boasting strict safety guarantees that are otherwise lost when candidate CBF-based control becomes infeasible.
The rest of the paper is organized as follows: Section 2 provides formal definitions for control invariance, CBFs, and DSMs; Section 3 proves that DSMs are CBFs, shows the relative degree of the constructed CBF is irrelevant for safety, and formulates the CBF-based optimization problem, which is a quadratic program (QP) for polyhedral input constraints; Section 4 presents the control-sharing property of CBFs and shows that DSMs have it; Section 5 provides a systematic approach based on Lyapunov functions to construct DSMs and highlights some desirable properties that such DSMs have; Section 6 presents the numerical simulations and comparisons; Section 7 concludes the paper.
@@ 1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\to\infty}\alpha(r)=\infty$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$, by ${\overline{\mathcal{S}}}$ the closure of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}={\overline{\mathcal{S}}}\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
@@ 2 Preliminaries
This section recalls the notion of control invariance, control barrier functions, and dynamic safety margins. Let $\mathbb{X}\subset\mathbb{R}^{n}$ and $\mathbb{U}\subset\mathbb{R}^{m}$ be the state and input spaces, respectively, and consider a control-affine system
@@
where $f:\mathbb{X}\to\mathbb{R}^{n}$ and $g:\mathbb{X}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions. The system is subject to state $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and input $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ constraints, where $\mathcal{X}\subset\mathbb{X}$ and $\mathcal{U}\subset\mathbb{U}$ are closed sets.
@@ 2.1 Control Invariance
Definition 1
[25] A compact set $\mathcal{C}\subset\mathbb{X}$ is control invariant if, $\forall\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$, there exists an admissible input signal $\boldsymbol{\mathbf{u}}:[0,\infty)\to\mathcal{U}$ such that, if they exist, at least one of the trajectories of the closed-loop system (1), with $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$, satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t$ in its interval of existence $\mathcal{J}(\boldsymbol{\mathbf{x}}_{0})\subset[0,\infty)$.
Note that this concept is sometimes called weak control invariance [25] because we don’t guarantee the existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s Theorem characterizing control invariance in terms of Bouligand’s tangent cone.
Theorem 1 (Nagumo)
[26] A compact set $\mathcal{C}\subset\mathbb{X}$ is control invariant if and only if
@@
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [25, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the compact set $\mathcal{C}$ is smooth, the tangent cone in the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2
[27, Proposition 4.3.7] Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 00 and suppose $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{X}$ is compact. The set $\mathcal{C}$ is control invariant if and only if
@@ 2.2 Control Barrier Functions
Definition 2
[6] A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
@@
$\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{X}.$ (5)
@@
Assume $\mathcal{U}\subset\mathbb{U}$ is compact. Let $\mathcal{C}\subset\mathbb{X}$, defined as in (5), be compact. Moreover, let 00 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
@@
Although the proof is based on [6], they showed this result for barrier functions, not control barrier functions. Let $\mathcal{C}\subset\mathbb{X}$ be control invariant and define the set $\mathcal{C}_{c}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:0\leq h(\boldsymbol{\mathbf{x}})\leq c\}$. Note that for any $c\geq 0$, $\mathcal{C}_{c}\subset\mathcal{C}$ is compact. Next, define the function $\hat{\alpha}:[0,\infty)\to\mathbb{R}$
$\hat{\alpha}(c)=-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{C}_{c}}\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right].$ (6)
By compactness of $\mathcal{U}$, the function $\xi:\mathcal{D}\to\mathbb{R}$ defined as
@@
is continuous on $\mathcal{D}$, making $\hat{\alpha}$ also continuous and non-decreasing. Furthermore, $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}})\big{)}$. To see why, let $\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$ be given and define $c_{0}=h(\boldsymbol{\mathbf{x}}_{0})$. There exists $\boldsymbol{\mathbf{u}}_{0}\in\mathcal{U}$ such that
$L_{f}h(\boldsymbol{\mathbf{x}}_{0})+L_{g}h(\boldsymbol{\mathbf{x}}_{0})\boldsymbol{\mathbf{u}}_{0}=\xi(\boldsymbol{\mathbf{x}}_{0})\geq\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{C}_{c_{0}}}\xi(\boldsymbol{\mathbf{x}})=-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}}_{0})\big{)}$
By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}_{0}$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
$\hat{\alpha}(0)=-\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}}\xi(\boldsymbol{\mathbf{x}})\leq 0$ (8)
There always exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$, yielding that $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$,
$\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right]\geq-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}.$
@@
Notice compactness of $\mathcal{U}$ is necessary for continuity of the function $\xi:\mathcal{D}\to\mathbb{R}$, and therefore of $\hat{\alpha}:[0,\infty)\to\mathbb{R}$. Without it, we lose necessity in Lemma 1. However, the sufficiency in Lemma 1 (i.e., the CBF certifies control invariance of its compact superlevel set) is maintained even when the compactness of $\mathcal{U}$ is exchanged for the less restrictive condition
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\mathcal{K}(\boldsymbol{\mathbf{x}})\neq\emptyset,$ (9)
@@
$\mathcal{K}(\boldsymbol{\mathbf{x}})=\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\geq 0\}.$ (10)

### v3

@@ Abstract
This paper presents an approach to design control barrier functions (CBFs) for arbitrary state and input constraints using tools from the reference governor literature. In particular, it is shown that dynamic safety margins (DSMs) are CBFs for an augmented system obtained by concatenating the state with a virtual reference. The proposed approach is agnostic to the relative degree and can handle multiple state and input constraints using the control-sharing property of CBFs. The construction of CBFs using Lyapunov-based DSMs is then investigated in further detail. Numerical simulations show that the method outperforms existing DSM-based approaches, while also guaranteeing safety and persistent feasibility of the associated optimization program.
@@ 1 Introduction
Given a state constraint set with a smooth boundary, it is usually straightforward to find a differentiable function of the state vector that is positive for constraint-admissible states, and negative otherwise. We call such functions candidate CBFs. They are commonly used in practice to design CBF-based controllers and, with tuning, can achieve great performance (e.g., [2]). However, candidate CBFs lack many safety guarantees that valid CBFs provide. In particular, the CBF-based optimization program governing the control law may become infeasible when using candidate CBFs, which is unacceptable for safety-critical systems. Ultimately, CBFs are certificates of control invariance. Therefore, if the original state constraint set is not control invariant, it is impossible to find a valid CBF associated to it. In these cases, we must resort to finding CBFs for a, hopefully large, control invariant subset of the state constraint set. In this work, we use tools from the reference governor literature [4] to systematically design valid CBFs given arbitrary state and input constraints.
@@ 1.1 Related Works
Most relevant to the present manuscript, we review some works that study the design of CBFs and feasibility of the CBF-based optimization program. In [14], sum of squares (SOS) techniques are used to synthesize CBFs for semi-algebraic sets and polynomial systems using an alternating descent algorithm. In [15], the authors use SOS techniques to synthesize polynomial control Lyapunov functions, and then present an extension to construct CBFs. In [16], the authors introduce a variable modifying the decay rate of the CBF to guarantee feasibility of the CBF-based program when the state is in the interior of the safe set even under input constraints. However, their approach provides no feasibility guarantee when the state is on the boundary of the safe set. In [17], the authors use control Lyapunov functions (CLFs) to design a continuum of valid CBFs and use tools from hybrid systems theory to study the resulting safety properties. In [18], the authors design CBFs for Euler–Lagrange systems. While the results are promising and the class of systems is relevant, the approach is limited to box constraints. In [19], the authors learn CBFs from human demonstrations but they provide no guarantees that the learned CBFs are valid. In [20], given a valid CBF, the authors use learning to construct a CBF for a larger control invariant set using a combination of safe and unsafe trajectories of the system. Nonetheless, finding the initial, valid CBF remains an open question. In [21], the authors use a backup control policy to enlarge a small (but easy to find) control invariant set. The enlarged control invariant set is then used to derive a backup CBF. These backup CBFs rarely have a closed form, which makes them difficult to implement. In [22], the authors use maximal output admissible sets (MOASs) to design discrete-time CBFs for arbitrary state and input constraints. While the approach works well for linear systems, finding the MOAS for nonlinear systems remains an open question.
@@
This work expands the underlying theme of [21, 22], where CBFs are obtained starting from a prestabilizing (or backup) controller. To do this, we adopt the notion of dynamic safety margin (DSM) from the explicit reference governor (ERG) framework [23]. Given a fixed reference, a DSM measures the distance to constraint violation of a prestabilized system trajectory as it converges to the associated equilibrium point. In [24], the authors define governor-parameterized barrier functions (PBFs), which are based on CBFs, and show that they are DSMs. While the results are limited to linear systems, their work is one of the few connections between the reference governor and CBF literatures. In contrast to their result, we show that DSMs are CBFs for an augmented system consisting of the concatenated state and reference vectors. These works serve as motivation for exploring the connections between reference governor approaches like the command governor (CG) [25], which filters a nominal reference, and CBF-based control, which filters a nominal input. Both approaches ultimately modify a nominal, but potentially unsafe signal to guarantee closed-loop safety (see Fig. 1).
@@ 1.2 Contributions
  • We leverage the control-sharing property presented in [9] to define a natural generalization of control barrier functions (CBFs) to (vector-valued) CBFs.
  • We show that dynamic safety margins (DSMs) are (vector-valued) CBFs for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM.
  • We show that the DSM-CBFs support constraints with arbitrary relative degree.
  • We formulate the DSM-CBF optimization-based policy and show that it is safe and persistently feasible. Moreover, the optimization problem is a quadratic program (QP) when the input constraints are polyhedral.
  • We provide conditions that guarantee the DSM-CBF optimization-based policy is locally Lipschitz continuous. This guarantees existence and uniqueness of the closed-loop trajectory.
  • We extend the Lyapunov-based DSMs presented in [23] to inadmissible references. This extension is needed to use them as DSM-CBFs. Moreover, we show that any class $\mathscr{K}_{\infty}$ function can be used with Lyapunov-based DSM-CBFs.
  • We present a comparison of our proposed approach with the explicit reference governor (ERG) [23], candidate CBFs and backup CBFs [21] in two nonlinear systems. The numerical simulations show that the DSM-CBFs we propose outperform the ERG, while boasting strict safety guarantees that are otherwise lost when candidate CBF-based control becomes infeasible. The examples also show that, while backup CBFs can achieve better performance than DSM-CBFs, their implementation is challenging in the presence of multiple constraints or loss of control authority.
The rest of the paper is organized as follows: Section 2 provides formal definitions for control invariance, CBFs, and DSMs; Section 3 proves that DSMs are CBFs, shows the relative degree of the constructed CBF is irrelevant for safety, formulates the CBF-based optimization problem, which is a quadratic program (QP) for polyhedral input constraints, and details some conditions that guarantee the DSM-CBF policy is locally Lipschitz continuous; Section 4 provides a systematic approach based on Lyapunov functions to construct DSMs and highlights some desirable properties that such DSMs have; Section 5 summarizes the presented CBF synthesis approach and contextualizes it with respect to the existing reference governor literature; Section 6 presents the numerical simulations and comparisons; Section 7 concludes the paper.
@@ 1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\to\infty}\alpha(r)=\infty$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}=\text{closure}(\mathcal{S})\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Given a subset $\mathcal{C}\subset\mathbb{R}^{n}$ of the state space $\mathbb{R}^{n}$, we generally denote by $\tilde{\mathcal{C}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ its counterpart in the augmented state-reference space. Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h(\boldsymbol{\mathbf{x}})/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
@@ 2 Preliminaries
This section recalls the notion of control invariance, control barrier functions, and dynamic safety margins. Consider a control-affine system
@@
where $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions. The system is subject to state $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and input $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ constraints, where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{U}\subset\mathbb{R}^{m}$ are closed sets.
@@ 2.1 Control Invariance
Definition 1 ([26])
A set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if, $\forall\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$, there exists an admissible input signal $\boldsymbol{\mathbf{u}}:[0,\infty)\to\mathcal{U}$ such that at least one of the trajectories of the closed-loop system (1), with $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$, satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t$ in its interval of existence $\mathcal{J}(\boldsymbol{\mathbf{x}}_{0})\subset[0,\infty)$.
Note that this concept is sometimes called weak control invariance [26] because it doesn’t require existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s theorem characterizing control invariance in terms of Bouligand’s tangent cone.
Theorem 1 (Nagumo [27])
A compact set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if and only if
@@
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [26, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the compact set $\mathcal{C}$ has a smooth boundary (i.e., can be represented as the superlevel set of a continuously differentiable function), the tangent cone on the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2 ([28, Proposition 4.3.7])
Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 00 and suppose $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$ is compact. The set $\mathcal{C}$ is control invariant if and only if
@@
Note that [28, Proposition 4.3.7] is more general than Theorem 2: it allows the set $\mathcal{C}$ to be defined using multiple continuously differentiable functions. We use this fact in upcoming results.
@@ 2.2 Control Barrier Functions
Definition 2 ([6])
A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
@@
$\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}.$ (5)
@@
Assume $\mathcal{U}\!\subset\!\mathbb{R}^{m}$ is compact. Let $\mathcal{C}\!\subset\!\mathbb{R}^{n}$, defined as in (5), be compact and let 00 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
@@
Although the proof is based on [6], they showed this result for barrier functions, not control barrier functions. Let $\mathcal{C}\subset\mathbb{R}^{n}$ be control invariant and define the set-valued map $C(c)=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid 0\leq h(\boldsymbol{\mathbf{x}})\leq c\}$. Note that for any $c\geq 0$, $C(c)\subset\mathcal{C}$ is compact in $\mathbb{R}^{n}$. Next, define the function $\hat{\alpha}:[0,\infty)\to\mathbb{R}$
$\hat{\alpha}(c)=-\inf_{\boldsymbol{\mathbf{x}}\in C(c)}\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right].$ (6)
Since $\mathcal{U}$ is compact, it follows by the extreme value theorem that $\xi:\mathcal{D}\to\mathbb{R}$, defined as
@@
is continuous on $\mathcal{D}$. Moreover, it can be shown that the set-valued map $C$ is both uppper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [29], that the value function $\hat{\alpha}$ is continuous on $[0,\infty)$. By construction, $\hat{\alpha}$ is non-decreasing. By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}(0)$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
$\hat{\alpha}(0)=-\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}}\xi(\boldsymbol{\mathbf{x}})\leq 0.$ (8)
There always exists a class $\mathscr{K}_{\infty}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$. Finally, let $\boldsymbol{\mathbf{x}}\in\mathcal{C}$ be given and note
$\xi(\boldsymbol{\mathbf{x}})\geq\inf_{\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}(h(\boldsymbol{\mathbf{x}}))}\xi(\boldsymbol{\mathbf{x}}_{0})=-\hat{\alpha}\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\geq-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}.$
So, we conclude that $h$ is a CBF.
@@
Notice compactness of $\mathcal{U}$ implies $\xi:\mathcal{D}\to\mathbb{R}$ is continuous, and therefore $\hat{\alpha}:[0,\infty)\to\mathbb{R}$. Without it, we lose necessity in Lemma 1. However, the sufficiency in Lemma 1 (i.e., the CBF certifies control invariance of its compact superlevel set) is maintained even when the compactness of $\mathcal{U}$ is exchanged for the less restrictive condition
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})\neq\emptyset,$ (9)
@@
$\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})\!=\!\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}\!\geq\!0\}.$ (10)